SkinnyMinnie20:

If m⦟PSM= 22°, find m⦟RSM and m⦟SPQ

Ultrilliam:

@Angle cough cough

SkinnyMinnie20:

1 attachment
Angle:

@SkinnyMinnie20 Are we assuming that this is a parallelogram?

SkinnyMinnie20:

yes

SkinnyMinnie20:

@Angle are you still here

Angle:

Yeah, I'm just trying to work it out; it's been a while since I've taken geometry

SkinnyMinnie20:

okay thank you

Angle:

Are you given any other information? like, is it a specific type of parallelogram? were there any other given angles before this question?

SkinnyMinnie20:

that's all I was given

Angle:

do they want numerical answers? or something like m<SPQ = m<RSM + m<PSM ?

SkinnyMinnie20:

I have no idea

Angle:

do you have an answer box?

Angle:

or is this a worksheet?

SkinnyMinnie20:

its a worksheet

Angle:

I see, then perhaps saying something like m<SPQ = m<RSM + m<PSM might work 1V7uUBksRI_Znd7qqtAO2A.png

Angle:

I'll break my picture down for you...

SkinnyMinnie20:

that's what I was thinking but what about the other two angles

Angle:

<PSM = <RQM by alternate interior angles so <RQM = 22 degrees and <SPQ = <QRS then <RQM + <QRS + <RSM = 180 because it is a triangle so 22 + <QRS + <RSM = 180 22 + <SPQ + <RSM = 180

SkinnyMinnie20:

im confused

Angle:

aye, the letters are confusing I'm sorry it's basically the picture I labeled (written out) + saying that 1 + 2 + 3 = 180 XwKfhRuwQPqRsNBZbO5K8Q.png

SkinnyMinnie20:

okay, so how would i find <RSM and <SPQ?

Angle:

(1) = 180 - (2) - (3) we know that angle (3) = 22 so angle (1) = 180 - 22 - (2)

Angle:

you would be stuck with saying <RSM = 158 - <SPQ

SkinnyMinnie20:

okay